# Low Photon Number Non-Invasive Imaging Through Time-Varying Diffusers

### v1

@@ 4 Discussion
We successfully demonstrated imaging through a dynamic diffuser at signal levels low enough to detect single photons. Our method, like techniques presented in [7, 6, 10], does not require physical alteration of the sample, nor does it necessitate characterizing the medium’s transmission matrix or focusing light through it. However, our approach is specifically optimized for low-photon conditions and dynamic diffusers – a scenario highly relevant for deep tissue imaging. This approach can be implemented without adaptive optics or wavefront shaping systems.
@@
We tested our method using a standard fluorescence microscope and applied it to various fluorescent samples. Notably, we successfully imaged astrocyte cells stained with the widely used dye Alexa Fluor 568, as well as various challenging objects constructed from fluorescent micro-spheres, further underscoring the versatility and effectiveness of our approach. Although we demonstrated our technique in microscopy, it has potential applications in various other fields of imaging. For instance, it can be utilized in LIDAR to cope with time-varying fog [22, 23], in astronomy for imaging distant, dim objects through a fluctuating atmosphere [24], and in endoscopy using multimode fibers [25, 26], where bends generate changing scattering scenarios.
@@
Single camera frames in our experiment contain low signals, reaching $0.33$ photons per pixel per frame (around $10^{6}$ per frame). These experiments are feasible thanks to recent advances in the development of single photon cameras. The camera that we utilized has a relatively low frame rate, capturing fewer than 5 frames per second for the full field of view required by the speckle size. A potential improvement in the data acquisition rate our setup is the use of SPAD arrays, which are increasingly being applied in imaging [27, 28, 29, 30, 31]. Recent advancements in this field have led to the development of megapixel SPAD arrays [29], making their practical application in widefield microscopy feasible. Cameras based on SPAD technology allow for single photon detection with high temporal resolution, providing a powerful tool for enhancing various bioimaging techniques and other applications. In our application, orders of magnitude improvement of temporal resolution offered by SPAD arrays will enable imaging through extremely fast varying diffusers.
@@ 5.1 Experimental setup
We designed and constructed our experimental setup, which consisted of a fluorescence wide-field microscope. For the objective, we used Nikon S Plan Fluor ELWD 20x/0.45, and for the excitation beam, we utilized a Spectra-Physics Millennia PRO laser with a wavelength of 532 nm. We utilized several neutral density filters to attenuate the laser beam. The fluorescence sample was placed within the working distance of the primary objective, while a rotating diffuser (Edmund Holographic Diffuser of 5° Diffusing Angle) was positioned adjacent to the microscope objective. To select the desired field of view in our experiment, we illuminated the sample from a transmissive side using another objective (Nikon Nikon S Plan Fluor ELWD 40x/0.60), which was usually out of focus to provide a large excitation spot.
To filter the fluorescence light from the excitation source, we employed a dichroic mirror (Edmund 552), a long-pass filter (Thorlabs FELH0550), and a notch filter Semrock NF01-532U-25. Additionally, to ensure good visibility of speckles, a band-pass filter (Thorlabs FBH590-10) was used. The tube lens in our setup was a Thorlabs TTL200-A with a focal length of 200 mm. The ORCA-Quest 2 qCMOS camera from Hamamatsu was positioned at the focal point of the tube lens.
@@
For more details of the experimental setup, please refer to Supplementary information.
@@ 5.2 Data acquisition
Initially, the computer sent a command to the Arduino to illuminate the sample and then to the camera to begin capturing light. The microcontroller ensured that the fluorescent sample was illuminated only during the camera’s light detection period to limit the photo-bleaching of the sample. It was removing the excitation laser beam immediately after the camera sent a trigger indicating the end of the frame, ensuring precise synchronization.
@@ 5.3 Data analysis
We performed several filtering techniques in our data processing to retrieve the object.
First, we divided by a 2D Gaussian function. The speckle image on the camera is not equally distributed across the sample, exhibiting a Gaussian envelope. By dividing by a 2D Gaussian function, we corrected for this uneven distribution.
Next, we applied a 2D Hann filter to smooth the edges of each frame, which helps mitigate edge effects in the Fourier domain.
After applying the Hann filter, we computed each frame’s 2D Fourier transform and calculated its modulus. We then averaged with the root mean square these transformed frames and convolved the resulting signal with a narrow 2D filter to remove speckle artifacts.
Finally, we applied the phase retrieval algorithm (the combination of hybrid input-output (HIO) and the error-reduction (ER) algorithms) to the prepared Fourier amplitude, to obtain a reconstructed object.
@@ 5.4 Sample preparation
Fluorescent microsphere samples were prepared using a drop-casting method. A concentrated solution of fluorescent microspheres (PS-FluoRed-Particles; 4.99 µm/SD=0.16 µm; 530/607 nm) was significantly diluted with ethanol. The solution was subjected to a 15-minute sonication bath to prevent the clustering of the fluorescent microspheres. A 50 µL drop of this diluted solution was then placed onto a microscope coverslip and allowed to air dry.
@@
Three-week-old primary mixed neuronal cultures on glass slides were fixed in PBS with 4% paraformaldehyde and 4% sucrose for 8 minutes at $37^{o}$. After washing three times with PBS containing 4% sucrose, cells were permeabilized with 0.25% Triton X-100 for 2 minutes. Blocking was performed in PBS with 3% BSA and 0.1% Triton X-100 for 1 hour at room temperature. Cultures were incubated overnight at $4^{o}$ with an anti-GFAP primary antibody (1:1000, #ab7260), followed by three PBS washes. Secondary antibody incubation was conducted for 1 hour at room temperature using Alexa Fluor 568-conjugated anti-rabbit antibody (1:500, #A-11011). After three PBS washes, samples were mounted on glass slides using Fluoromount-G solution (#00-4958-02).
@@
Acknowledgements We would like to express our sincere gratitude to Diana Legutko for preparing the Alexa Fluor 568-stained astrocyte samples, and to Anat Daniel for insightful discussions and comments about the experimental setup. We also acknowledge the support of the following funding agencies: Narodowe Centrum Nauki (grants 2023/49/N/ST7/04195 and 2022/47/B/ST7/03465); Fundacja na rzecz Nauki Polskiej (FIRST TEAM project POIR.04.04.00-00-3004/17-00); the European Regional Development Fund (POIR.04.04.00-00-3004/17-00); HORIZON EUROPE Marie Skłodowska-Curie Actions (FLORIN ID 101086142); and the French Government Scholarship for Ph.D. Cotutelle/Codirection.

### v2

@@ 4 Discussion
We successfully demonstrated imaging through a dynamic diffuser at signal levels low enough to detect single photons, conditions under which no imaging has been demonstrated so far. This scenario is highly relevant for deep-tissue imaging, where scattering and low photon counts are significant challenges. Our approach does not require physical alteration of the sample, nor does it necessitate characterizing the medium’s transmission matrix or focusing light through it. Our approach does not require adaptive optics or wavefront shaping systems. Our method is robust, remaining unaffected by spatially uncorrelated noise.
Moreover, our method is resilient to fluorescence bleaching and minor sample movements during multi-frame acquisition. Although bleaching reduces signal intensity, the spatial distribution of fluorophores in densely stained samples remains intact, preserving essential spatial information despite reduced Fourier amplitudes. By averaging the Fourier transform moduli across frames, we accurately recover the object’s Fourier transform. Similarly, since translational motion only introduces phase in the Fourier domain, our modulus-based approach naturally compensates for such shifts—only significant rotations or structural changes would impact the reconstruction. Our method works also in the case of static diffusers (see Supplementary Information Section 2.1.6).
In Section 2 of the Supplementary Information, we provide a detailed discussion of the theoretical foundations of our approach. We analyze the required number of frames for a given photon count per frame and demonstrate that our approach has no strict lower bound on signal intensity, allowing object reconstruction by accumulating a sufficient number of frames. In addition, we explore the spatial resolution of our imaging technique and its signal-to-noise ratio.
@@
We validated our approach using a standard fluorescence microscope and applied it to various fluorescent samples. Notably, we successfully imaged astrocyte cells stained with the widely used dye Alexa Fluor 568, as well as various challenging objects constructed from fluorescent micro-spheres, further underscoring the versatility and effectiveness of our approach. Our approach has potential applications in various other fields of imaging. For instance, it can be utilized in LIDAR to cope with time-varying fog [22, 23], in astronomy for imaging distant, dim objects through a fluctuating atmosphere [24], and in endoscopy using multimode fibers [25, 26], where bends generate changing scattering scenarios.
Our approach has no lower bound on fluorescence intensity and can operate even with just 2 photons per frame. However, achieving performance with as few as 2 photons per frame may require a substantial number of frames, as detailed in Section 2.1.6 of the Supplementary Information. Single camera frames in our experiment contain low signals, reaching $0.14$ photons per pixel per frame (around $3\times 10^{5}$ per 9 Mpx frame). These experiments are feasible thanks to recent advances in the development of single photon cameras. The camera that we utilized has a relatively low frame rate, capturing fewer than 5 frames per second for the full field of view required by the speckle size. Since our experiment requires thousands of camera frames, the total acquisition time for a single image reconstruction is typically a couple of minutes. Nonetheless, this acquisition duration is well-suited for many bioimaging applications, such as embryonic development, cell division, and synaptic modifications.
@@
A potential improvement in the data acquisition rate our setup is the use of SPAD arrays, which are increasingly being applied in imaging [27, 28, 29, 30, 31]. Recent advancements in this field have led to the development of megapixel SPAD arrays [29], making their practical application in widefield microscopy feasible. Cameras based on SPAD technology allow for single photon detection with high temporal resolution, providing a powerful tool for enhancing various bioimaging techniques and other applications. Some SPAD arrays offer sub-picosecond resolution for photon arrival time tagging [29]. By correlating detected photons within a fixed time window while the diffuser remains stationary, we can optimize data utilization and significantly enhance acquisition efficiency, enabling much faster fluorescence imaging. In our case, orders of magnitude improvement of temporal resolution offered by SPAD arrays will enable imaging through extremely fast varying diffusers, making it suitable for more realistic scenarios.
@@ 5.1 Experimental setup
We designed and constructed our experimental setup, which consisted of a fluorescence wide-field microscope. For the objective, we used Nikon S Plan Fluor ELWD 20x/0.45, and for the excitation beam, we utilized a Spectra-Physics Millennia PRO laser with a wavelength of 532 nm. We utilized several neutral density filters to attenuate the laser beam. The fluorescence sample was placed within the working distance of the primary objective, while a rotating diffuser (Edmund Holographic Diffuser of 5° Diffusing Angle) was positioned adjacent to the microscope objective (between the objective and the sample). To select the desired field of view in our experiment, we illuminated the sample from a transmissive side using another objective (Nikon S Plan Fluor ELWD 40x/0.60), which was usually out of focus to provide an adjustable excitation spot.
To separate the fluorescence light from the excitation source, we employed a dichroic mirror (Edmund 552), a long-pass filter (Thorlabs FELH0550), and a notch filter Semrock NF01-532U-25. Additionally, to ensure good visibility of speckles, a band-pass filter (Thorlabs FBH590-10) was used. The tube lens in our setup was a Thorlabs TTL200-A with a focal length of 200 mm. The ORCA-Quest 2 qCMOS camera from Hamamatsu was positioned at the focal point of the tube lens.
@@
For more details of the experimental setup, please refer to the Supplementary Information Section 1.
@@ 5.2 Data acquisition
Initially, the computer sent a command to the Arduino to illuminate the sample and then to the camera to begin the acquisition. The microcontroller ensured that the fluorescent sample was illuminated only during the camera’s light detection period to limit the photo-bleaching of the sample. It was removing the excitation laser beam immediately after the camera sent a trigger indicating the end of the frame.
@@ 5.3 Data analysis
We performed several filtering techniques in our data processing to retrieve the object. The speckle image on the camera is not equally distributed across the sample, exhibiting a Gaussian envelope. By dividing by a 2D Gaussian function, we corrected for this uneven distribution. Next, we applied a 2D Hann filter to smooth the edges of each frame, which helps mitigate edge effects in the Fourier domain. After applying the Hann filter, we computed each frame’s 2D Fourier transform and calculated its modulus. We then averaged these transformed frames with the root mean square and convolved the resulting signal with a narrow 2D filter to remove speckle artifacts. Finally, we applied the phase retrieval algorithm (the combination of hybrid input-output (HIO) and the error-reduction (ER) algorithms) to the prepared Fourier amplitude, to obtain a reconstructed object.
@@ 5.4 Sample preparation
Fluorescent microsphere samples were prepared using a drop-casting method. A concentrated solution of fluorescent microspheres (PS-FluoRed-Particles; 4.99 µm/SD=0.16 µm; 530/607 nm) was significantly diluted with ethanol. The solution was subjected to a 15-minute sonication bath to prevent clustering of the fluorescent microspheres. A 50 µL drop of this diluted solution was then placed onto a microscope coverslip and allowed to air dry.
@@
Three-week-old primary mixed neuronal cultures on glass slides were fixed in PBS with 4% paraformaldehyde and 4% sucrose for 8 minutes at $37^{o}$C. After washing three times with PBS containing 4% sucrose, cells were permeabilized with 0.25% Triton X-100 for 2 minutes. Blocking was performed in PBS with 3% BSA and 0.1% Triton X-100 for 1 hour at room temperature. Cultures were incubated overnight at $4^{o}$C with an anti-GFAP primary antibody (1:1000, #ab7260), followed by three PBS washes. Secondary antibody incubation was conducted for 1 hour at room temperature using Alexa Fluor 568-conjugated anti-rabbit antibody (1:500, #A-11011). After three PBS washes, samples were mounted on glass slides using Fluoromount-G solution (#00-4958-02).
@@
Data availability The data used in this study is available from the corresponding author upon reasonable request.
[…]
Acknowledgements We would like to express our sincere gratitude to Diana Legutko for preparing the Alexa Fluor 568-stained astrocyte samples, and to Anat Daniel for insightful discussions and comments about the experimental setup. We want to acknowledge the support of the following funding agencies: Foundation for Polish Science (FIRST TEAM project FENG.02.02-IP.05-0253/23); Narodowe Centrum Nauki (grants 2023/49/N/ST7/04195 and 2022/47/B/ST7/03465); National Centre for Research and Development QuantERA II projects: QM3 (QuantERAII/02/QM3/03/2024) and EXTRASENS (QuantERAII/02/EXTRASENS/02/2024); HORIZON EUROPE Marie Skłodowska-Curie Actions (FLORIN ID 101086142); and the French Government Scholarship for Ph.D. Cotutelle/Codirection.‘
[…]
Conflict of interest The authors declare no competing interests.
[…]
Contributions A.M. contributed to conceiving the project idea together with S.G. and R.L., designed and built the experimental setup, prepared the fluorescence microsphere samples, conducted the experiments, performed all simulations, developed and investigated the data analysis methods, analyzed the data, prepared all figures, and wrote the manuscript. W.Z. contributed to the theoretical framework, supported improvements in the data analysis, and authored the supplementary section on theoretical investigations. P.S. provided key input on constructing the experimental platform, contributed to data acquisition and theory, and wrote the supplementary section detailing the experimental setup. B.G. supported the theoretical investigations and co-authored relevant parts of the theoretical supplement. S.G. contributed to its initial conception and provided critical insights and guidance throughout the project. R.L. conceived the project, coordinated its development, and provided essential guidance and input across all stages. All authors reviewed and revised the manuscript and supplementary materials.